OpenStudy (anonymous):

Lagrange help: I have the answers but can someone check it for me? f(x,y)=x^2y^2 9x^2+4y^2=36

OpenStudy (anonymous):

fx(x,y) = 2xy^2 fy(x,y) = 2x^2y g(x,y) = 9x^2+4y^2 -36 gx(x,y) = 18x gy(x,y) = 8y

OpenStudy (anonymous):

\[2xy ^{2}=\lambda18x\] \[2x ^{2}y=\lambda8y\]

OpenStudy (anonymous):

\[\frac{2xy ^{2}}{18x}, x \neq0\]

OpenStudy (anonymous):

and 2x^2y/8y

OpenStudy (anonymous):

my question is i dont know if (0,0) or (x,0), (0,y) is a solution

OpenStudy (anonymous):

i have y = +-sqrt(4.5) and x = +-sqrt(2)

OpenStudy (anonymous):

then i fill that in in f(x,y)

OpenStudy (anonymous):

but i dont know if zero is also a solution :S

hero (hero):

Okay, what was the question?

hero (hero):

Oh

hero (hero):

I can't exactly say that I can really help here